OpenStudy (anonymous):

For f(x)=log2x and g(x)=4^x, find a) g(f(2)) b) f(g^(-1)(64))

OpenStudy (solomonzelman):

first find the g( f(x) ) by plugging in what the f(x) is equal to, INTO the g(x), INSTEAD of x.

OpenStudy (anonymous):

So \[\log _{2}4=2\]?

OpenStudy (solomonzelman):

f( g(x) ) = 4 ^ ( log_2 x) like this

OpenStudy (solomonzelman):

then plug in 2 for x/

OpenStudy (solomonzelman):

I mean this is g( f(x) ) g of f, not f of g... sorry.

OpenStudy (solomonzelman):

the right side of what I wrote is correct.

OpenStudy (anonymous):

not to butt in but since is this is going to be a number, you can compute the number \[f(2)=\log_2(2)\] (which should be more or less obvious) and then compute \(g\) of the result

OpenStudy (solomonzelman):

\(\Large\color{black}{\color{red}{f(x)=\log_2x} }\) \(\Large\color{black}{\color{blue}{g(x)=4^x} }\) \(\Large\color{black}{\color{blue}{g(\color{red}{f(x)})=4^{\color{red}{\log_2x}}} }\)

OpenStudy (solomonzelman):

\(\Large\color{black}{\color{blue}{g(\color{red}{f(x)})=4^{\color{red}{\log_2x}}} }\) \(\Large\color{black}{\color{blue}{g(\color{red}{f(\color{darkgoldenrod}{2})})=4^{\color{red}{\log_2(\color{darkgoldenrod}{2})}}} }\)

OpenStudy (solomonzelman):

then take it from there..

OpenStudy (anonymous):

Ohhh okay, I get it now, I was forming the equation completely wrong, I don't know why I was doing it the way I was but thank you! It is much appreciated.

OpenStudy (solomonzelman):

anytime:)